# Motion-Aware Transformer for Multi-Object Tracking

### v2

@@ 4.5 Comparison on SportsMOT
We evaluate MATR on the SportsMOT [cui2023sportsmot] dataset, Table 4 presents comparisons between MATR and state-of-the-art methods on the SportsMOT test set, all under settings without additional datasets for pretraining or fine-tuning.
Detection performance on SportsMOT is generally strong across methods, as indicated by MOTA scores consistently above 90. However, association accuracy (AssA) remains comparatively limited, typically around or below 60. Since SportsMOT already provides a sufficient amount of training data, we emphasize improving association accuracy rather than relying on external datasets. It is also important to note that SportsMOT annotations cover only athletes and exclude spectators, which differs significantly from datasets such as CrowdHuman [shao2018crowdhuman]. While traditional tracking-by-detection methods can effectively filter irrelevant detections, these annotation inconsistencies introduce additional challenges for end-to-end approaches.
Thanks to the baseline results reported by [gao2023memotr], we are able to perform fair comparisons without external data. MATR significantly improves upon previous methods across all evaluation metrics. In particular, MATR achieves a new state-of-the-art HOTA score of 72.7, representing gains of 2.7 points over MeMOTR [gao2023memotr] and 4.6 points over OC-SORT [cao2023observation]. While our detection accuracy (DetA) improves only slightly compared to OC-SORT (+0.5 points), the most notable gains are observed in association metrics: +2.9 AssA and +3.2 IDF1 compared to MeMOTR. These results highlight the effectiveness of MATR’s motion-aware strategy, which consistently optimizes both tracking and detection queries, leading to substantial improvements in association performance while preserving strong detection accuracy.
@@ 4.6 Comparison on BDD100k
As the largest and most diverse MOT dataset, BDD100k [bdd100k] poses substantial challenges for tracking methods. Recently, TETer [li2022tracking] introduced the TETA metric, a comprehensive evaluation framework designed specifically for BDD100k, along with their method. Although TETer achieves the highest localization accuracy (mLocA = 47.2), like other tracking-by-detection methods, it struggles with association metrics, which reduces its overall performance on mTETA and mHOTA.
@@
As shown in Table 5, MATR achieves state-of-the-art performance on nearly all metrics. Compared to MOTR, MATR provides large improvements across the board (+4.0 mTETA, +4.6 mHOTA, +6.0 mLocA, +8.0 mAssocA, and +5.7 mAssA), demonstrating the effectiveness of our MAT module over the original QIM design. For prior SOTA methods such as CO-MOT [yan2023bridging] and MeMOTR [gao2023memotr], which rely on detection-specific tricks or historical frame information, MATR also achieves superior results. Although CO-MOT and MeMOTR both reach similar association scores (56.2 and 56.7 mAssocA, respectively), our method improves by about 3 absolute points, establishing a new state of the art at 59.0 mAssocA.

### v3

@@ 4.5 Comparison on SportsMOT
We evaluate MATR on the SportsMOT [8] dataset, Table 4 presents comparisons between MATR and state-of-the-art methods on the SportsMOT test set, all under settings without additional datasets for pretraining or fine-tuning.
Detection performance on SportsMOT is generally strong across methods, as indicated by MOTA scores consistently above 90. However, association accuracy (AssA) remains comparatively limited, typically around or below 60. Since SportsMOT already provides a sufficient amount of training data, we emphasize improving association accuracy rather than relying on external datasets. It is also important to note that SportsMOT annotations cover only athletes and exclude spectators, which differs significantly from datasets such as CrowdHuman [18]. While traditional tracking-by-detection methods can effectively filter irrelevant detections, these annotation inconsistencies introduce additional challenges for end-to-end approaches.
Thanks to the baseline results reported by [9], we are able to perform fair comparisons without external data. MATR significantly improves upon previous methods across all evaluation metrics. In particular, MATR achieves a new state-of-the-art HOTA score of 72.7, representing gains of 2.7 points over MeMOTR [9] and 4.6 points over OC-SORT [5]. While our detection accuracy (DetA) improves only slightly compared to OC-SORT (+0.5 points), the most notable gains are observed in association metrics: +2.9 AssA and +3.2 IDF1 compared to MeMOTR. These results highlight the effectiveness of MATR’s motion-aware strategy, which consistently optimizes both tracking and detection queries, leading to substantial improvements in association performance while preserving strong detection accuracy.
@@ 4.6 Comparison on BDD100k
As the largest and most diverse MOT dataset, BDD100k [29] poses substantial challenges for tracking methods. Recently, TETer [12] introduced the TETA metric, a comprehensive evaluation framework designed specifically for BDD100k, along with their method. Although TETer achieves the highest localization accuracy (mLocA = 47.2), like other tracking-by-detection methods, it struggles with association metrics, which reduces its overall performance on mTETA and mHOTA.
@@
As shown in Table 5, MATR achieves state-of-the-art performance on nearly all metrics. Compared to MOTR, MATR provides large improvements across the board (+4.0 mTETA, +4.6 mHOTA, +6.0 mLocA, +8.0 mAssocA, and +5.7 mAssA), demonstrating the effectiveness of our MAT module over the original QIM design. For prior SOTA methods such as CO-MOT [26] and MeMOTR [9], which rely on detection-specific tricks or historical frame information, MATR also achieves superior results. Although CO-MOT and MeMOTR both reach similar association scores (56.2 and 56.7 mAssocA, respectively), our method improves by about 3 absolute points, establishing a new state of the art at 59.0 mAssocA.